Hello. Ok the problem that I am having with my homework is this. The question says to solve each system of equations by using the substitution method. I am already aware that the problem here will have two answers so here it is
y=2x
x+y=12
I have tried to subtract the y value from the x+y=12, but I am not sure what I am doing!

Answer by checkley71(8403)

IF Y=2X & X+Y=12 THEN X+2X=12 OR 3X=12 THUS X=12/3 OR X=4 & Y=2X OR Y=8.
X+Y=12 SO 4+8=12.